14.

Complete the first step in fraction multiplication by choosing the fraction that represents the unsimplified product of-and The simplified product of these two fractions is Complete the first step in determining the sum of and The final, simplified sum of these two fractions is Complete the first step in dividing the fraction by the fraction The final, simplified quotient of these two fractions is
Answer #1

14. 1) (6x2)/(7x3) = 12/21

Simplified product = 4/7

2) 1/4 + 5/8

Taking LCM of denominator, we get 8

= (2 + 5)/8 = 7/8

Simplified sum is 7/8.

3) 6/7 / 3/2

6/7 x 2/3 = (6 x 2)/(7 x 3) = 12/21

Simplified quotient = 12/21 = 4/7